OpenStudy (anonymous):

graph the line with slope -4 and passing through (-2,4)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-2\quad ,&4) \end{array} \\\quad \\ slope = m= -4 \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"}\)

OpenStudy (isaiah.feynman):